Find the inverse of the function
f(x)=4+ \sqrt{x-2} .
State the domains and ranges of both the function and the inverse function in terms of intervals of real numbers.
obtain the graph of f , its inverse, and g(x)=x in the same system of axes. About what pair (a, a) are (11, 7) and (7, 11) reflected about?
1
Expert's answer
2020-10-01T15:40:36-0400
"Solution"

To find the inverse of a function there are 2 steps:

  1. Everywhere there's a "y, f(x), g(x)," etc, swap it with x
  2. Solve for "y" in the rewritten function

I'm going to assume that the function is "f(x) = 4 + \sqrt{x-2}" where the "x-2" is under the radical sign


So the first step is to swap the "f(x)" and "x" . The new function would be

"x = 4 + \sqrt{f(x) -2}"

Now we'll solve for "f(x)" in the new function

"x = 4 + \sqrt{f(x) -2}\\

x-4 = \sqrt{f(x) -2}\\

(x-4)^2 = f(x) -2\\

x^2 - 8x + 16 = f(x) -2\\

x^2 - 8x + 18 = f(x)"

So the inverse function, which we write as "f^{-1}(x)" is


"f^{-1}(x) = x^2 - 8x + 18"


Domain of "f(x)"


"x-2 \geq 0\\
\implies x \geq 2"

Range of "f(x)"

"[4, \infin)"

From the graph, the points "(11, 7)" and "(7, 11)" reflected about point "(9,9)".